# SparseSwaps: Tractable LLM Pruning Mask Refinement at Scale

### v1

@@ Abstract
The resource requirements of Neural Networks can be significantly reduced through pruning – the removal of seemingly less important parameters. However, with the rise of Large Language Models (LLMs), full retraining to recover pruning-induced performance degradation is often prohibitive and classical approaches such as global magnitude pruning are suboptimal on Transformer architectures. State-of-the-art methods hence solve a layer-wise mask selection problem, the problem of finding a pruning mask which minimizes the per-layer pruning error on a small set of calibration data. Exactly solving this problem to optimality using Integer Programming (IP) solvers is computationally infeasible due to its combinatorial nature and the size of the search space, and existing approaches therefore rely on approximations or heuristics. In this work, we demonstrate that the mask selection problem can be made drastically more tractable at LLM scale. To that end, we decouple the rows by enforcing equal sparsity levels per row. This allows us to derive optimal 1-swaps (exchanging one kept and one pruned weight) that can be computed efficiently using the Gram matrix of the calibration data. Using these observations, we propose a tractable and simple 1-swap algorithm that warm starts from any pruning mask, runs efficiently on GPUs at LLM scale, and is essentially hyperparameter-free. We demonstrate that our approach reduces per-layer pruning error by up to 60% over Wanda (Sun et al., 2023) and consistently improves perplexity and zero-shot accuracy across state-of-the-art GPT architectures.
@@ 1 Introduction
Pruning after training (Han et al., 2015; Gale et al., 2019; Lin et al., 2020; Hoefler et al., 2021; Zimmer et al., 2025) is a state-of-the-art technique to reduce the resource requirements of neural networks. A simple yet effective approach to obtain such sparse models starts from a pretrained dense model, removes seemingly unimportant parameters based on their magnitude, and requires retraining to compensate for pruning-induced performance degradation. However, while the inexpensive, data-free magnitude criterion has often achieved strong performance on traditional architectures (Gale et al., 2019; Zimmer et al., 2023b), pruning has undergone a paradigm shift with the rise of large pretrained foundation models, particularly LLMs.
First, the size of the models has shifted the focus toward retraining-free pruning criteria, as retraining is often computationally expensive if not infeasible, with parameter-efficient fine-tuning (Lialin et al., 2023; Zimmer et al., 2023a) being an exception. Secondly, systematic activation outliers (Dettmers et al., 2022) and highly important super-weights (Yu et al., 2025) in sufficiently large Transformers (Vaswani et al., 2017) have rendered magnitude pruning no better than random pruning for LLMs (Sun et al., 2023; Yin et al., 2023). Lastly, state-of-the-art methods (Frantar & Alistarh, 2023; Sun et al., 2023; Zhang et al., 2024) prune layer-wise: they split the pruning problem into per-layer subproblems, pruning layers sequentially and independently using a small calibration dataset to estimate parameter importance. Rather than optimizing the global loss, such approaches minimize a per-layer local pruning loss. Specifically, for a single layer with calibration input matrix $X\in\mathbb{R}^{d_{in}\times B}$ and weights $W\in\mathbb{R}^{d_{out}\times d_{in}}$, the objective becomes
@@
Solving this combinatorial mask selection problem to optimality is NP-hard due to feature correlations: selecting $k$ of $d_{out}\cdot d_{in}$ weights yields a cardinality-constrained binary quadratic program (a best-subset selection variant). Even for a single row $i$, the problem reduces to
$\min_{m_{i}}{\left\|w_{i}^{\top}X-(m_{i}\odot w_{i})^{\top}X\right\|}_{F}^{2}=\min_{m_{i}}\sum_{k=1}^{B}\left(\sum_{j=1}^{d_{in}}(1-m_{ij})w_{ij}X_{jk}\right)^{2},$
where $w_{i}\in\mathbb{R}^{d_{in}}$ and $m_{i}\in{\{0,1\}}^{d_{in}}$ denote the $i$-th row of $W$ and $M$, respectively. While IP solvers could theoretically provide optimal solutions, the combinatorial search over mask entries makes this infeasible for LLMs. In practice, existing methods therefore relax Equation 1 or approximate it.
However, with deployed LLMs now serving millions of users, it becomes increasingly worthwhile to invest substantial resources to obtain pruned models that reach high performance, because the pruning cost is paid once during training whereas inference costs scale with the number of requests. In this work, we revisit the per-layer mask selection problem and demonstrate that it can be operationalized at LLM scale, enabling monotone improvements with each optimization step rather than relying on proxy importance scores. To that end, we observe that enforcing equal sparsity-level across rows ensures row-wise separability that yields independent objectives. This makes the problem drastically more tractable and leads to good practical performance for LLMs. Instead of trying to obtain exact solutions via IP solvers, we instead propose a GPU-accelerated local optimization algorithm based on 1-swaps (exchanging one kept and one pruned weight) that perform exact and efficient local refinement with incremental cost updates using the Gram matrix $G=XX^{\top}$ to monotonically decrease the objective from any warm start.
The resulting method, which we term SparseSwaps, can start from any warm-start mask, evaluates the exact per-row quadratic loss, and is scalable, parallelizable across rows, almost hyperparameter-free, and deterministic for a fixed warm start. With only few 1-swap iterations, it can reduce the per-layer pruning error by up to 60% compared to Wanda and improves final perplexity and zero-shot accuracy across architectures. Our approach is a post-hoc refinement of existing pruning methods that can significantly improve upon the state of the art for unstructured, per-row, or $N\text{:}M$ sparsity.
@@
    SparseSwaps: a practical post-hoc pruning algorithm. Building on these observations, we propose SparseSwaps, a plug-and-play 1-swap refinement that starts from any warm-start mask and monotonically decreases the exact per-row objective under per-row or $N\text{:}M$ constraints. In particular, SparseSwaps is almost hyperparameter-free, completely parallelizable across rows and scalable to LLMs.
@@
    Computational study. We verify our hypotheses on state-of-the-art Generative Pretrained Transformer (GPT) architectures and demonstrate that SparseSwaps delivers large reductions in local pruning error (up to 60% per-layer error reduction over Wanda) and strong perplexity and zero-shot gains across a wide range of different LLMs. We conduct a series of ablations highlighting the advantages and drawbacks of the proposed approach.
@@
Post-training pruning has a long history, and while magnitude pruning (Janowsky, 1989; Han et al., 2015) is among the most popular criteria, it is not the only one (cf. LeCun et al., 1989; Hassibi & Stork, 1993; Molchanov et al., 2016; Yeom et al., 2019); see Hoefler et al. (2021) for a comprehensive review. Despite their simplicity, magnitude-based methods have been shown to produce sparse models competitive with far more complex algorithms for convolutional architectures (Gale et al., 2019; Zimmer et al., 2023b). For LLMs, however, magnitude pruning is argued to be unsuitable (Yin et al., 2023). Consequently, there is growing interest in criteria beyond magnitude that achieve high performance on LLMs, and do so without requiring an expensive retraining procedure (Kwon et al., 2022; Frantar & Alistarh, 2023; Sun et al., 2023; Zhang et al., 2024). In this work, we develop a post-hoc refinement of existing methods, rather than proposing a new criterion. A related approach, DSnoT (Zhang et al., 2023), also performs iterative weight swaps but differs significantly in its optimization strategy. Inspired by dynamic sparse training (cf. Evci et al., 2020), DSnoT prunes and regrows weights based on expected reconstruction-error improvements, using feature means and variances as surrogates. While effective, it does not guarantee a monotonic decrease in the true pruning error, whereas our method does. We compare the two empirically and find that SparseSwaps consistently outperforms DSnoT.
@@ 2 Methodology
In the following, we use uppercase letters for matrices ($W$, $X$, $M$) and lowercase letters for scalars and vectors. Matrix entries are denoted $W_{ij}$ for the element in row $i$, column $j$. Rows of matrices are denoted with lowercase subscripts: $w_{i}$ represents the $i$-th row of matrix $W$. Row and column slices use colon notation: $X_{j,:}$ for the $j$-th row and $X_{:,k}$ for the $k$-th column. We use $\odot$ for element-wise multiplication, ${\left\|\cdot\right\|}_{F}$ for Frobenius norm, and ${\left\|\cdot\right\|}_{2}$ for $\ell_{2}$ norm.
@@ 2.1 Preliminaries
2.1.1 Equal sparsity-level across rows does not need to be detrimental
@@
${\left\|WX-(M\odot W)X\right\|}_{F}^{2}=\sum_{i=1}^{d_{out}}{\left\|w_{i}^{\top}X-(m_{i}\odot w_{i})^{\top}X\right\|}_{2}^{2}$
where $w_{i}\in\mathbb{R}^{d_{in}}$ and $m_{i}\in{\{0,1\}}^{d_{in}}$ denote the $i$-th row of $W$ and $M$, respectively. This alone does not make the corresponding minimization problem row-separable under unstructured sparsity, since the matrix cardinality constraint couples rows. In contrast, semi-structured patterns like per-row sparsity (keep $k$ per row) or $N\text{:}M$ (prune MNM\!-\!N per block of $M$ weights) enforce equal per-row sparsity, meaning that the rows are fully decoupled by definition. We therefore focus on the decoupled case, allowing to treat each row separately and reducing the problem to
$\min_{m_{i}}{\left\|w_{i}^{\top}X-(m_{i}\odot w_{i})^{\top}X\right\|}_{F}^{2}=\min_{m_{i}}\sum_{k=1}^{B}\left(\sum_{j=1}^{d_{in}}(1-m_{ij})w_{ij}X_{jk}\right)^{2}$ (2)
for each row $i\in\{1,\ldots,d_{out}\}$. Note that, for LLMs, Sun et al. (2023) observe that row-wise sparsity benefits performance for both Wanda and magnitude pruning. We therefore argue that enforcing per-row sparsity rather than unstructured sparsity is justified and need not harm final performance, at least for LLMs.
@@ 2.1.2 Avoiding intermediate value caching through the Gram matrix formulation
$L\stackrel{{\scriptstyle\mathrm{\scriptscriptstyle def}}}{{=}}{\left\|w^{\top}X-(m\odot w)^{\top}X\right\|}_{F}^{2}={\left\|(w-m\odot w)^{\top}X\right\|}_{F}^{2}=(w-m\odot w)^{\top}G(w-m\odot w).$
Hence, the loss depends on $X$ only through the Gram matrix $G$, which can be accumulated on-the-fly as calibration samples pass through the layer: $G=\sum_{b=1}^{B}X_{:,b}X_{:,b}^{\top}$. Unlike the per-row formulation in the introduction, which would require caching all $B\cdot d_{in}$ intermediate products $w_{j}X_{jk}$, we only need to maintain the $d_{in}\times d_{in}$ matrix $G$, which is a reduction from $\mathcal{O}(B\cdot d_{in})$ to $\mathcal{O}(d_{in}^{2})$ since $d_{in}$ is typically much smaller than $B$.
Remark 1.
A different (but in practice slightly less efficient) perspective on this reduction is through the unitary invariance of the Frobenius norm used in our pruning objective: for any matrix $A$ and unitary matrix $U$ (i.e., $U^{-1}=U^{\top}$), we have ${\left\|AU\right\|}_{F}={\left\|A\right\|}_{F}$. This property enables significant computational savings through Singular Value Decomposition (SVD) compression. Precisely, let $X=U\Sigma V^{\top}$ be the SVD of calibration data $X\in\mathbb{R}^{d_{in}\times B}$. Since $B>d_{in}$, we can write $\Sigma=[\Sigma^{\prime}\mid 0]$ with $\Sigma^{\prime}\in\mathbb{R}^{d_{in}\times d_{in}}$ containing the singular values on its diagonal. The compressed representation is simply $X^{\prime}=U\Sigma^{\prime}\in\mathbb{R}^{d_{in}\times d_{in}}$. Letting $w_{p}=w-m\odot w$ for brevity, the key insight is that pruning decisions remain equivalent under this compression:
${\left\|w_{p}X\right\|}_{F}^{2}={\left\|w_{p}U\Sigma V^{\top}\right\|}_{F}^{2}={\left\|w_{p}U\Sigma\right\|}_{F}^{2}={\left\|w_{p}U[\Sigma^{\prime}\mid 0]\right\|}_{F}^{2}={\left\|w_{p}U\Sigma^{\prime}\right\|}_{F}^{2}={\left\|w_{p}X^{\prime}\right\|}_{F}^{2},$
@@ 2.1.3 Efficient 1-swap evaluation through efficient cost lookups and updates
where we define the reconstruction residual $r\stackrel{{\scriptstyle\mathrm{\scriptscriptstyle def}}}{{=}}\sum_{j\in\mathcal{P}}w_{j}\phi_{j}\in\mathbb{R}^{B}$, the total contribution of all pruned weights to the layer output. Hence, clearly, the loss is $L={\left\|r\right\|}_{2}^{2}=r^{\top}r$.
We define the correlation vector $c=(c_{1},\ldots,c_{d_{in}})^{\top}\in\mathbb{R}^{d_{in}}$ with entries
$c_{i}\stackrel{{\scriptstyle\mathrm{\scriptscriptstyle def}}}{{=}}{\langle\phi_{i},r\rangle}={\langle\phi_{i},\sum_{j\in\mathcal{P}}w_{j}\phi_{j}\rangle}=\sum_{j\in\mathcal{P}}w_{j}{\langle\phi_{i},\phi_{j}\rangle}=\sum_{j\in\mathcal{P}}w_{j}G_{ij},$
@@
$\Delta L_{u,p}$ $={\left\|r^{\prime}\right\|}_{2}^{2}-{\left\|r\right\|}_{2}^{2}={\left\|r+w_{u}\phi_{u}-w_{p}\phi_{p}\right\|}_{2}^{2}-{\left\|r\right\|}_{2}^{2}$
$=2w_{u}{\langle\phi_{u},r\rangle}+w_{u}^{2}{\left\|\phi_{u}\right\|}_{2}^{2}-2w_{p}{\langle\phi_{p},r\rangle}+w_{p}^{2}{\left\|\phi_{p}\right\|}_{2}^{2}-2w_{u}w_{p}{\langle\phi_{u},\phi_{p}\rangle}.$
Using $c_{i}={\langle\phi_{i},r\rangle}$ and $G_{ij}={\langle\phi_{i},\phi_{j}\rangle}$, this simplifies to
$\Delta L_{u,p}=2w_{u}c_{u}+w_{u}^{2}G_{uu}-2w_{p}c_{p}+w_{p}^{2}G_{pp}-2w_{u}w_{p}G_{up}.$ (3)
Given the precomputed Gram matrix $G$ and correlation vector $c$, each swap evaluation requires only scalar lookups. Evaluating all possible swaps therefore costs $\mathcal{O}(|\mathcal{U}|\cdot|\mathcal{P}|)$ total. By systematically testing all $((d_{in}-|\mathcal{P}|)\cdot|\mathcal{P}|)$ possible 1-swap operations (adding one of $|\mathcal{U}|=d_{in}-|\mathcal{P}|$ unpruned weights to $\mathcal{P}$, removing one of $|\mathcal{P}|$ pruned weights from $\mathcal{P}$) evaluating the improvement using the above expression, we iteratively pick a best swap and update the mask until we have reached a satisfactory solution or one optimal w.r.t. 1-swap operations. The only issue that remains is to update the correlation vector after each swap.

### v2

@@ Abstract
The resource requirements of neural networks can be significantly reduced through pruning – the removal of seemingly less important parameters. However, for LLMs, full retraining to recover pruning-induced performance degradation is often prohibitive and classical approaches such as magnitude pruning are suboptimal on Transformers. State-of-the-art methods hence solve a layer-wise mask selection problem: finding a pruning mask that minimizes per-layer pruning error on a small set of calibration data. Exactly solving this problem is computationally infeasible due to its combinatorial nature and the size of the search space, and existing approaches rely on approximations or heuristics. We demonstrate that the mask selection problem can be made drastically more tractable at LLM scale. To that end, we decouple the rows by enforcing equal sparsity levels per row. This allows us to derive optimal 1-swaps (exchanging one kept and one pruned weight) computable efficiently via the Gram matrix. We propose a simple 1-swap algorithm that warmstarts from any pruning mask, runs efficiently on GPUs at LLM scale, and is essentially hyperparameter-free. Our approach reduces per-layer pruning error by up to 60% over Wanda (Sun et al., 2024) and consistently improves perplexity and zero-shot accuracy across state-of-the-art GPT architectures.
Machine Learning, ICML
@@ 1 Introduction
Pruning after training (Han et al., 2015; Gale et al., 2019; Lin et al., 2020; Hoefler et al., 2021; Zimmer et al., 2025) is a state-of-the-art technique to reduce the resource requirements of neural networks. A simple yet effective approach to obtain such sparse models starts from a pretrained dense model, removes seemingly unimportant parameters based on their magnitude, and requires retraining to compensate for pruning-induced performance degradation. However, while the inexpensive, data-free magnitude criterion has often achieved strong performance on traditional architectures (Gale et al., 2019; Zimmer et al., 2023b), pruning has undergone a paradigm shift with the rise of pretrained foundation models such as Large Language Models (LLMs).
First, the size of the models has shifted the focus toward retraining-free pruning criteria, as retraining is often computationally expensive if not infeasible, with parameter-efficient fine-tuning (Lialin et al., 2023; Zimmer et al., 2023a) being an exception. Secondly, systematic activation outliers (Dettmers et al., 2022) and highly important super-weights (Yu et al., 2025) in sufficiently large Transformers (Vaswani et al., 2017) have rendered magnitude pruning no better than random pruning for LLMs (Sun et al., 2024; Yin et al., 2023). Lastly, state-of-the-art methods (Frantar & Alistarh, 2023; Sun et al., 2024; Zhang et al., 2024a) prune layer-wise: they split the pruning problem into per-layer subproblems, pruning layers sequentially and independently using a small calibration dataset to estimate parameter importance. Rather than optimizing the global loss, such approaches minimize a per-layer local pruning loss. Specifically, for a single layer with calibration input matrix $X\in\mathbb{R}^{d_{in}\times B}$ and weights $W\in\mathbb{R}^{d_{out}\times d_{in}}$, the objective becomes
@@
Solving this combinatorial mask selection problem to optimality is NP-hard due to feature correlations: selecting $k$ of $d_{out}\cdot d_{in}$ weights yields a cardinality-constrained binary quadratic program (a best-subset selection variant). Even for a single row $i$ the problem remains hard, despite reducing to
$\min_{m_{i}}{\left\|w_{i}^{\top}X-(m_{i}\odot w_{i})^{\top}X\right\|}_{2}^{2}$
where $w_{i}\in\mathbb{R}^{d_{in}}$ and $m_{i}\in{\{0,1\}}^{d_{in}}$ denote the $i$-th row of $W$ and $M$, respectively. While Integer Programming (IP) solvers could theoretically provide optimal solutions, the combinatorial search over mask entries makes this infeasible for LLMs. In practice, existing methods therefore relax Equation 1 or approximate it.
However, with deployed LLMs now serving millions of users, it becomes increasingly worthwhile to invest substantial resources to obtain pruned models that reach high performance, because the pruning cost is paid once during training whereas inference costs scale with the number of requests. In this work, we revisit the per-layer mask selection problem and demonstrate that it can be operationalized at LLM scale, enabling monotone improvements with each optimization step rather than relying on proxy importance scores. To that end, we observe that enforcing equal sparsity levels across rows ensures row-wise separability that yields independent objectives. This makes the problem drastically more tractable and still leads to good practical performance for LLMs. Instead of trying to obtain exact solutions via IP solvers, we instead propose a GPU-accelerated local optimization algorithm based on 1-swaps (exchanging one kept and one pruned weight) that perform exact and efficient local refinement with incremental cost updates using the Gram matrix $G=XX^{\top}$ to monotonically decrease the objective from any warmstart.
Figure 1: Per-layer relative reduction in local pruning error compared to Wanda. The plot shows results for LLaMA-3.1-8B, 60% unstructured sparsity and 100 1-swap iterations.
The resulting method, which we term SparseSwaps, can start from any warmstart mask, evaluates the exact per-row quadratic loss, and is scalable, parallelizable across rows, almost hyperparameter-free, and deterministic for a fixed warmstart. With only a few 1-swap iterations, it can reduce the per-layer pruning error by up to 60% compared to Wanda and improves final perplexity and zero-shot accuracy across architectures. Our approach is a post-hoc refinement of existing pruning methods that can significantly improve upon the state of the art for unstructured, per-row, or $N\text{:}M$ sparsity.
@@
    SparseSwaps: a practical post-hoc pruning algorithm. Building on these observations, we propose SparseSwaps, a plug-and-play 1-swap refinement that starts from any warmstart mask and monotonically decreases the exact per-row objective under per-row or $N\text{:}M$ constraints. In particular, SparseSwaps is almost hyperparameter-free, completely parallelizable across rows and scalable to LLMs.
@@
    Computational study. We verify our hypotheses on state-of-the-art Generative Pretrained Transformer (GPT) architectures and demonstrate that SparseSwaps delivers large reductions in local pruning error (up to 60% per-layer error reduction over Wanda) and strong perplexity and zero-shot gains across a wide range of different LLMs. Surprisingly, when pruning-induced degradation is low, we find that while SparseSwaps still consistently and drastically reduces the local error, these gains do not necessarily translate to perplexity improvements, highlighting inherent limitations of layer-wise mask optimization as an approximation to the global objective. We conduct a series of ablations highlighting the advantages and drawbacks of the proposed approach.
@@
Post-training pruning has a long history, and while magnitude pruning (Janowsky, 1989; Han et al., 2015) is among the most popular criteria, it is not the only one (cf. LeCun et al., 1989; Hassibi & Stork, 1993; Molchanov et al., 2016; Yeom et al., 2019); see Hoefler et al. (2021) for a comprehensive review. Despite their simplicity, magnitude-based methods have been shown to produce sparse models competitive with far more complex algorithms for convolutional architectures (Gale et al., 2019; Zimmer et al., 2023b). For LLMs, however, magnitude pruning is argued to be unsuited (Yin et al., 2023). Consequently, there is growing interest in criteria beyond magnitude that achieve high performance on LLMs, and do so without requiring an expensive retraining procedure (Kwon et al., 2022; Frantar & Alistarh, 2023; Sun et al., 2024; Zhang et al., 2024a). In this work, we develop a post-hoc refinement of existing methods, rather than proposing a new criterion. A related approach, DSnoT (Zhang et al., 2024b), also performs iterative weight swaps but differs significantly in its optimization strategy. Inspired by dynamic sparse training (cf. Evci et al., 2020), DSnoT prunes and regrows weights based on expected reconstruction-error improvements, using feature means and variances as surrogates. While effective, it does not guarantee a monotonic decrease in the true pruning error, whereas our method does. We compare the two empirically and find that SparseSwaps consistently outperforms DSnoT.
@@ 2 Methodology
In the following, we use uppercase letters for matrices ($W$, $X$, $M$) and lowercase letters for scalars and vectors. Matrix entries are denoted $W_{ij}$ for the element in row $i$, column $j$. Rows of matrices are denoted with lowercase subscripts: $w_{i}$ represents the $i$-th row of matrix $W$. Row and column slices use colon notation: $X_{j,:}$ for the $j$-th row and $X_{:,k}$ for the $k$-th column. We use $\odot$ for element-wise multiplication, ${\left\|\cdot\right\|}_{F}$ for the Frobenius norm, and ${\left\|\cdot\right\|}_{2}$ for the $\ell_{2}$ norm.
@@ 2.1 Preliminaries
2.1.1 Row Decoupling through Equal Per-Row Sparsity
@@
${\left\|(W-M\odot W)X\right\|}_{F}^{2}=\sum_{i=1}^{d_{out}}{\left\|w_{i}^{\top}X-(m_{i}\odot w_{i})^{\top}X\right\|}_{2}^{2}$
where $w_{i}\in\mathbb{R}^{d_{in}}$ and $m_{i}\in{\{0,1\}}^{d_{in}}$ denote the $i$-th row of $W$ and $M$, respectively. If we enforce equal sparsity levels per row, the rows become fully decoupled and can be treated independently, reducing the problem to
$\min_{m_{i}}{\left\|w_{i}^{\top}X-(m_{i}\odot w_{i})^{\top}X\right\|}_{2}^{2}$ (2)
for each row $i\in\{1,\ldots,d_{out}\}$. This is naturally the case for per-row sparsity constraints (keep $k$ weights per row) and semi-structured $N\text{:}M$ patterns (prune MNM\!-\!N per block of $M$ weights). For unstructured sparsity, however, the global cardinality constraint ${\left\|M\right\|}_{0}\leq k$ couples rows. Sun et al. (2024) observe that enforcing row-wise sparsity benefits performance for both Wanda and magnitude pruning on LLMs. We therefore assume per-row sparsity even in nominally unstructured settings, enabling row decoupling without necessarily harming final performance.
As a side note, we can now easily derive the Wanda criterion by upper bounding the error with Jensen’s inequality and noting that positive scaling preserves minima, i.e.,
$\min_{m_{i}}{\left\|w_{i}^{\top}X-(m_{i}\odot w_{i})^{\top}X\right\|}_{F}^{2}$
$=$ $\min_{m_{i}}\sum_{k=1}^{B}\left(\sum_{j=1}^{d_{in}}(1-m_{ij})w_{ij}X_{jk}\right)^{2}$ (3)
$\leq$ $\min_{m_{i}}\sum_{k=1}^{B}\left(\sum_{j=1}^{d_{in}}(1-m_{ij})^{2}w_{ij}^{2}X_{jk}^{2}\right)$
$=$ $\min_{m_{i}}\sum_{j=1}^{d_{in}}(1-m_{ij})^{2}w_{ij}^{2}{\left\|X_{j,:}\right\|}_{2}^{2}.$ (4)
Equation 4 is solved by pruning entries with the smallest saliency $|w_{ij}|\cdot{\left\|X_{j,:}\right\|}_{2}$, i.e., precisely the Wanda criterion. Thus, Wanda optimizes an upper bound to the original problem that ignores within-row interactions, making the combinatorial problem tractable.
@@ 2.1.2 Avoiding Intermediate Value Caching Through the Gram Matrix Formulation
$L$ $\stackrel{{\scriptstyle\mathrm{\scriptscriptstyle def}}}{{=}}{\left\|w^{\top}X-(m\odot w)^{\top}X\right\|}_{F}^{2}={\left\|(w-m\odot w)^{\top}X\right\|}_{F}^{2}$
$=(w-m\odot w)^{\top}G(w-m\odot w).$
Hence, the loss depends on $X$ only through the Gram matrix $G$, which can be accumulated on-the-fly as calibration samples pass through the layer: $G=\sum_{b=1}^{B}X_{:,b}X_{:,b}^{\top}$. Unlike the per-row formulation of Equation 3, which would require caching all $B\cdot d_{in}$ intermediate products $w_{j}X_{jk}$, we only need to maintain the $d_{in}\times d_{in}$ matrix $G$, which is a reduction from $\mathcal{O}(B\cdot d_{in})$ to $\mathcal{O}(d_{in}^{2})$ (with $d_{in}$ typically being much smaller than $B$).
@@ 2.1.3 Efficient 1-Swap Evaluation Through Efficient Cost Lookups and Updates
where we define the reconstruction residual
$r\stackrel{{\scriptstyle\mathrm{\scriptscriptstyle def}}}{{=}}\sum_{j\in\mathcal{P}}w_{j}\phi_{j}\in\mathbb{R}^{B},$
the total contribution of all pruned weights to the layer output. Hence, clearly, the loss is
$L={\left\|r\right\|}_{2}^{2}=r^{\top}r.$
We define the correlation vector $c\in\mathbb{R}^{d_{in}}$ with entries
$c_{i}\stackrel{{\scriptstyle\mathrm{\scriptscriptstyle def}}}{{=}}{\langle\phi_{i},r\rangle}={\langle\phi_{i},\sum_{j\in\mathcal{P}}w_{j}\phi_{j}\rangle}$ $=\sum_{j\in\mathcal{P}}w_{j}{\langle\phi_{i},\phi_{j}\rangle}$
$=\sum_{j\in\mathcal{P}}w_{j}G_{ij},$
@@
$\Delta L_{u,p}$ $\stackrel{{\scriptstyle\mathrm{\scriptscriptstyle def}}}{{=}}{\left\|r^{\prime}\right\|}_{2}^{2}-{\left\|r\right\|}_{2}^{2}={\left\|r+w_{u}\phi_{u}-w_{p}\phi_{p}\right\|}_{2}^{2}-{\left\|r\right\|}_{2}^{2}$
$=2w_{u}{\langle\phi_{u},r\rangle}+w_{u}^{2}{\left\|\phi_{u}\right\|}_{2}^{2}-2w_{p}{\langle\phi_{p},r\rangle}$
$\quad+w_{p}^{2}{\left\|\phi_{p}\right\|}_{2}^{2}-2w_{u}w_{p}{\langle\phi_{u},\phi_{p}\rangle}.$
Using $c_{i}={\langle\phi_{i},r\rangle}$ and that $G_{ij}={\langle\phi_{i},\phi_{j}\rangle}$, this simplifies to
$\Delta L_{u,p}$ $=2w_{u}c_{u}+w_{u}^{2}G_{uu}$
$\quad-2w_{p}c_{p}+w_{p}^{2}G_{pp}-2w_{u}w_{p}G_{up}.$ (5)
Given the precomputed Gram matrix $G$ and correlation vector $c$, each swap evaluation requires only scalar lookups. The theoretical complexity of evaluating all possible swaps therefore equals $\mathcal{O}(|\mathcal{U}|\cdot|\mathcal{P}|)$. By systematically testing all $((d_{in}-|\mathcal{P}|)\cdot|\mathcal{P}|)$ possible 1-swap operations (adding one of $|\mathcal{U}|=d_{in}-|\mathcal{P}|$ unpruned weights to $\mathcal{P}$, removing one of $|\mathcal{P}|$ pruned weights from $\mathcal{P}$) evaluating the improvement using the above expression, we iteratively pick a best swap and update the mask until we have reached a satisfactory solution or one optimal w.r.t. 1-swap operations. The only issue that remains is to update the correlation vector after each swap.